OpenStudy (anonymous):

If a 2.5 g sample of gold is hammered into an equilateral triangle .28mm thick, how long are the sides?

OpenStudy (anonymous):

I think I know what to do, but I would still like some extra help!

OpenStudy (anonymous):

Someone please help?

OpenStudy (ineedhelplz):

@FaiqRaees

OpenStudy (anonymous):

@Redcan do you think you could help me?

OpenStudy (faiqraees):

\[\large\rm Volume=\frac{base*height*0.28}{2} \]\[\large\rm Volume=mass~in~grams\]

OpenStudy (anonymous):

Really? It's that simple? Thank you!

OpenStudy (anonymous):

So the total volume would be the lenght of the sides? @FaiqRaees

OpenStudy (faiqraees):

the base will be length of side

OpenStudy (anonymous):

How would I find that?

OpenStudy (faiqraees):



OpenStudy (anonymous):

Um... Okay, let me figure this out.

OpenStudy (anonymous):

how would the 2.5g tie in with this?

OpenStudy (faiqraees):

just find the value of hieght in terms of x

OpenStudy (anonymous):

This is a density and volume related problem

OpenStudy (faiqraees):

no density isn't given

OpenStudy (anonymous):

So how could I tie in the 2.5g and the .28mm of thickness with this?

OpenStudy (faiqraees):

just please find out the value of height, let me take care of the rest

OpenStudy (anonymous):

Ok one second

OpenStudy (anonymous):

Wait, how exactly would I do that? I don't have any values for x

OpenStudy (faiqraees):

just in terms of x

OpenStudy (anonymous):

\[x^{2}\]

OpenStudy (anonymous):

I have to go, but I will post again later. Thanks for the help!

OpenStudy (faiqraees):

No the height will be \[\large\rm Hypotenuse^2=Perpendicular^2+Height^2 \]\[\large\rm Height^2=Hypotenuse^2-Perpendicular^2\]\[\large\rm Height^2=x^2-(x/2)^2\] \[\large\rm Height=\sqrt{x^2-(x/2)^2}\]

OpenStudy (faiqraees):

got it?

OpenStudy (faiqraees):

\[\large\rm Volume=\frac{base*height*0.28}{2} \]\[\large\rm 2.5=\frac{x*\sqrt{x^2-(x/2)^2}*0.28}{2} \]\[\large\rm 5=x*\sqrt{x^2-(x/2)^2}*0.28 \]\[\large\rm 17.85=x*\sqrt{x^2-(x/2)^2}\]\[\large\rm 17.85^2=(x*\sqrt{x^2-(x/2)^2})^2\]\[\large\rm 318.9=x^2*(x^2-(x/2)^2)\]\[\large\rm 318.9=x^2*(x^2-x^2/4)\]\[\large\rm 318.9=x^4-x^4/4\]\[\large\rm 318.9=3x^4/4\]\[\large\rm 425.5=x^4\]\[\large\rm 4.54=x\]

OpenStudy (faiqraees):

length of side = 4.54mm